OpenStudy (anonymous):

given the equation for a circle, identify its center and radius (x-2)^2+(y-3)^2=25

OpenStudy (anonymous):

center: (2,3) radius 5

OpenStudy (anonymous):

how did you get that? lol

OpenStudy (anonymous):

if the equation is (x-a)^2+(y-b)^2=c^2 then the center is (a,b) and the radius is c

OpenStudy (anonymous):

can i ask you another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

x^2+(y-7)^2=49

OpenStudy (anonymous):

that is (x-0)^2+(y-7)^2=7^2 so (0,7) and 7

OpenStudy (anonymous):

what about x^2=y^2=1 would it be (0,0) and 1?

OpenStudy (anonymous):

x^2+y^2=1 for this yes

OpenStudy (anonymous):

yay i got it right! lol thankyou

OpenStudy (anonymous):

your welcome